PLEASE HELP: I have to find the total amount due for each simple interest loan for $6100 at 5.7% interest for 4 years 9 months.
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

find the total amount due for each simple interest loan for $6100 at 5.7% interest for 4 years 9 months.
:
Find the interest due for 4.75 yrs
:
i = p*r*t
p = 6100
r = .057
t = 4.75
:
i = 6100 * .057 * 4.75
i = 1651.58
add to the principal
1651.58 + 6100 = $7,751.58

Answer by ikleyn(53937)

.
The total amount is


A = 6100*(1 + 0.057*4.75),


where  0.057 = 5.7%  and  4.75 = 4 years and 9 months.


Calculate the final result on your own.